OpenStudy (anonymous):

Steps for this integral, please: ∫[sin(x)]/[1+(cos(x))^2] dx =-arctan(cos(x))

OpenStudy (anonymous):

so you know the answer but want help with the stepS?

OpenStudy (turingtest):

huh? definite integral or indefinite?

OpenStudy (turingtest):

\[\int{\sin x\over1+\cos^2x}dx\]

OpenStudy (turingtest):

\[\int{\sin x\over1+\cos^2x}dx=-\tan^{-1}(\cos x)+C\]and we want to prove it, right?

OpenStudy (anonymous):

right

OpenStudy (turingtest):

\[u=\cos x\]\[du=-\sin xdx\]\[\int{\sin x\over1+\cos^2x}dx=-\int{du\over1+u^2}=-\tan^{-1}u+C\]\[=-\tan^{-1}(\cos x)+C\]do you know that last integral?

OpenStudy (anonymous):

Thank you for your time So arctan x=tan^-1x ?

OpenStudy (turingtest):

yes, just a different notation

OpenStudy (anonymous):

Thank you, once again for your help.

OpenStudy (turingtest):

anytime :D and if you don't know how to show that\[\int{dx\over a^2+x^2}=\frac1a\tan^{-1}(\frac xa)+C\]and you know how to do trig substitution integrals, I could prove that for you, but only if you wanted...

OpenStudy (anonymous):

Thanks, I'm good for now :)